OpenStudy (anonymous):

solve for x: logx+log(x+21)=2

OpenStudy (anonymous):

the answer is 4 I just don't know how to get there

myininaya (myininaya):

log(y)+log(z)=log(yz) tried to used that here

myininaya (myininaya):

log[x(x+21)]=2

OpenStudy (sasogeek):

alexa have you made an attempt to solve the question?

OpenStudy (anonymous):

yes i have and I come up with something completely different. im not sure whether to use the quotient, product, or power rule

OpenStudy (sasogeek):

combine the product and quotient rule

myininaya (myininaya):

to get rid of the log write base 10 on both sides so we have x(x+21)=10^2

OpenStudy (sasogeek):

log(x(x+21)=2 logx^2+21x=2 =>10^2=x^2+21

OpenStudy (sasogeek):

ummmm 21x

myininaya (myininaya):

100=x^2+21x

myininaya (myininaya):

right

OpenStudy (sasogeek):

simple quadratics

OpenStudy (sasogeek):

alexa I hope its alright now... problem solved?

OpenStudy (anonymous):

I can go from here, thanks

OpenStudy (sasogeek):

you're welcome :)

OpenStudy (sasogeek):

heyy btw alexa, you got msn or fb or some personal contact we can talk later, I'd appreciate it :/

myininaya (myininaya):

hey alex make sure you check your answers because we can't have log(negative)

OpenStudy (anonymous):

sorry but I dont like to give my personal information out unless I personally know the individual, no offense. and thanks, I will.

OpenStudy (sasogeek):

np, understood :)